# Enhancing Generative Auto-bidding with Offline Reward Evaluation and Policy Search

### v2

@@ A.1 RL-based Auto-bidding Methods
Auto-bidding plays a critical role in online advertising by automatically placing bids, allowing advertisers to participate efficiently in real-time auctions (Balseiro et al., 2021a; Deng et al., 2021; Balseiro et al., 2021b). The auto-bidding problem can be modeled as a Markov Decision Process and addressed using reinforcement learning techniques. USCB (He et al., 2021) proposes a unified solution to the constrained bidding problem, employing an RL method, DDPG (Silver et al., 2014), to dynamically adjust parameters to an optimal bidding strategy. Mou et al. (2022) design a sustainable online reinforcement learning framework that iteratively alternates between online explorations and offline training, alleviating the sim2rel problem. A few studies explore multi-agent RL for auto-bidding (Jin et al., 2018; Guan et al., 2021; Wen et al., 2022), while several focus on budget allocation and bidding strategies in multi-channel scenarios using RL-based approaches (Wang et al., 2023; Deng et al., 2023; Duan et al., 2025). Importantly, offline RL methods such as BCQ (Fujimoto et al., 2019), CQL (Kumar et al., 2020), IQL (Kostrikov et al., 2022), and MOPO (Yu et al., 2020) have demonstrated significant potential in this domain. These methods allow policy learning from pre-collected datasets without requiring online interaction. Moreover, offline RL, such as Diffusion-QL (Wang et al., 2022), adopts generative models as the policy model architecture for better expressive capacity.
@@ C.5 Proof of Theorem 3
$\leq\delta_{D}+(1+k)\sqrt{T}R_{m}\mathbb{E}_{y\sim p_{D}(y)}[W_{1}(p_{\theta^{*}}(\tau|y^{*}),p_{D}(\tau|y))],$ (41)
@@
$\mathbb{E}_{y\sim p_{D}(y)}[W_{1}(p_{\theta^{*}}(\tau|y^{*}),p_{D}(\tau|y))]\leq\sqrt{\frac{2}{\rho}\mathbb{E}_{y\sim p_{D}(y)}[D_{KL}(p_{\theta^{*}(\tau|y)}\|p_{D}(\tau|y))]}$ (42)
Let $\delta_{M}\triangleq\mathbb{E}_{y\sim p_{D}(y)}[D_{KL}(p_{\theta^{*}}(\tau|y)\|p_{D}(\tau|y))]$ be the distance between the optimal trajectory distribution and the offline dataset trajectory distribution. We have:
@@
$|J(\hat{\theta})-L(\hat{\theta})|\leq\delta_{D}+(1+k)\sqrt{T}R_{m}[(1+\epsilon)y_{m}L_{p}+\sqrt{2\delta_{K}/\rho}],$ (45)
Overall, combining the results in Eq. 43, Eq. C.5, and Eq. 45, we have:
@@ C.6 Proof of Score Gradient
Then, we have:
@@
Appendix D AIGB-Pearl Algorithm Summary
Algorithm 1 summarizes the training process of AIGB-Pearl. The development of AIGB-Pearl is supported by ROLL (Wang et al., 2025).
@@ Appendix D AIGB-Pearl Algorithm Summary
Input : Offline dataset $\mathcal{D}$, desired condition $y^{*}$, hyper-parameters $\beta_{1}$, $\beta_{2},\beta_{3}$.
@@
// Determining the Lipschitz Value
@@
Set the Lipschitz constraint value $L_{e}$ for the evaluator and $L_{p}$ for the planner to be bigger than the Lipschitz value of $y(\tau)$ and $p_{D}(\tau|y)$, respectively.
// Training the trajectory evaluator
@@
// Training the generative planner
@@
  Update $\theta$ by maximizing Eq. 3.2.2.
@@ E.3 Real-world Experiments on TargetROAS Bidding Problem
Table 7: Overall performance of TargetROAS in real-world A/B test, involving 300k advertisers over 22 days.
@@ E.4 Pathological Trajectory Behavior Explanation
E.5 Training Stability
We present additional comparisons between the training curves of the offline RL with bootstrapping and those of AIGB-Pearl in Fig. 6, Fig. 7, Fig. 8, and Fig. 9 concerning:
@@ E.5 Training Stability
We can see that the offline RL method tends to suffer from significant instability throughout training, showing high variance across different seeds. In contrast, AIGB-Pearl achieves much smoother and more consistent learning progress, demonstrating the improved training stability.

### v3

@@ A.1 RL-based Auto-bidding Methods
Auto-bidding plays a critical role in online advertising by automatically placing bids, allowing advertisers to participate efficiently in real-time auctions (Balseiro et al., 2021a; Deng et al., 2021; Balseiro et al., 2021b). The auto-bidding problem can be modeled as a Markov Decision Process and addressed using reinforcement learning techniques. USCB (He et al., 2021) proposes a unified solution to the constrained bidding problem, employing an RL method, DDPG (Silver et al., 2014), to dynamically adjust parameters to an optimal bidding strategy. Mou et al. (2022) design a sustainable online reinforcement learning framework that iteratively alternates between online explorations and offline training, alleviating the sim2rel problem. A few studies explore multi-agent RL for auto-bidding (Jin et al., 2018; Guan et al., 2021; Wen et al., 2022), while several focus on budget allocation and bidding strategies in multi-channel scenarios using RL-based approaches (Wang et al., 2023; Deng et al., 2023; Duan et al., 2025). Importantly, offline RL methods such as BCQ (Fujimoto et al., 2019), CQL (Kumar et al., 2020), IQL (Kostrikov et al., 2022), and MOPO (Yu et al., 2020) have demonstrated significant potential in this domain. These methods allow policy learning from pre-collected datasets without requiring online interaction. Moreover, offline RL, such as Diffusion-QL (Wang et al., 2022b), adopts generative models as the policy model architecture for better expressive capacity.
@@ C.5 Proof of Theorem 3
$\leq\delta_{D}+(1+k)\sqrt{T}R_{m}\mathbb{E}_{y\sim p_{D}(y)}\bigg[W_{1}(p_{\theta^{*}}(\tau|y^{*}),p_{D}(\tau|y))\bigg],$ (41)
@@
$\mathbb{E}_{y\sim p_{D}(y)}[W_{1}(p_{\theta^{*}}(\tau|y^{*}),p_{D}(\tau|y))]\leq\sqrt{\frac{2}{\rho}\mathbb{E}_{y\sim p_{D}(y)}[D_{KL}(p_{\theta^{*}}(\tau|y^{*})\|p_{D}(\tau|y))]}$ (42)
Let $\delta_{M}\triangleq\mathbb{E}_{y\sim p_{D}(y)}[D_{KL}(p_{\theta^{*}}(\tau|y^{*})\|p_{D}(\tau|y))]$ be the distance between the optimal trajectory distribution and the offline dataset trajectory distribution. We have:
@@
$|J(\hat{\theta})-L(\hat{\theta})|\leq\delta_{D}+(1+k)\sqrt{T}R_{m}\bigg[(1+\epsilon)y_{m}L_{p}+\sqrt{2\delta_{K}/\rho}\bigg],$ (45)
Overall, combining the results in Eq. C.5, Eq. 43, Eq. C.5, and Eq. 45, we have:
[…]
$\quad+k\sqrt{T}R_{m}\bigg[\sqrt{2\delta_{M}/\rho}+(1+\epsilon)y_{m}L_{p}+\sqrt{2\delta_{K}/\rho}\bigg]$
$\quad+\delta_{D}+(1+k)\sqrt{T}R_{m}\bigg[(1+\epsilon)y_{m}L_{p}+\sqrt{2\delta_{K}/\rho}\bigg]$
$=2\delta_{D}+(1+2k)\sqrt{T}R_{m}\bigg[\sqrt{2\delta_{M}/\rho}+\sqrt{2\delta_{K}/\rho}+(1+\epsilon)y_{m}L_{p}\bigg].$ (46)
@@ C.6 Proof of Score Gradient
Then, we can derive the gradient as follows:
@@
In conclusion, we have:
$\nabla_{\theta}L(\theta)=\mathbb{E}_{s_{1:T}\sim p_{\theta}(s_{1:T}|y^{*})}\bigg[\sum_{t}\nabla_{\theta}\log p_{\theta}(s_{t}|s_{1:t-1},y^{*})\hat{y}_{\phi}(\tau)\bigg].$ (49)
Appendix D AIGB-Pearl Practical Algorithm
D.1 Additional Designs For Evaluator Accuracy Enhancement
To further enhance the reliability of the trajectory evaluator, we design two additional techniques for the evaluator learning. Specifically, as described in the following, we (i) integrate an LLM embedding into its input feature for better representational capacity; and (ii) employ the pair-wise loss for better score estimation accuracy. Fig. 6 illustrates the complete evaluator learning losses. The effectiveness of these two methods is studied in Appendix E.6.
D.1.1 LLM Embedding Enhancement
Motivated by the success of integrating user-specific features into the recommendation system (Chen et al., 2015), we incorporate advertiser-specific features into the trajectory evaluator to enhance its representational capacity for better scoring accuracy. Note that some advertiser-specific features are textual in nature—such as product titles, categories, and reviews—and are therefore difficult to incorporate directly into the vectorized trajectory $\tau$. To address this, we construct a prompt containing all such textual attributes and employ a pre-trained large language model (LLM), QWen2.5-1.5B-Instruct, with general world knowledge to extract $T$ representation, which we refer to as the LLM embedding. Specifically, we use the output of the last hidden layer as the LLM embedding. This embedding is then used as an additional positional encoding in the evaluator. The prompt template is given below:
LLM Prompt Template. I am an [advertising platform] advertiser, operating the [brand name] brand in the category of [category name], classified as a [advertiser tier] tier advertiser. I have a product titled [product name] currently running in advertising campaigns. This product belongs to the leaf category of [leaf category], with a price of [product price], positioned in the [price range] price range. Its price ranks within the top [price ranking in the leaf category] % in the leaf category. Historical Average Performance: The product generates an average of [average daily transactions] daily transactions, with a GMV of [average daily GMV] driven by advertising. It receives an average [average daily impressions] daily impressions from search and recommendation traffic, [click numbers] clicks, [average daily BuyCnt] BuyCnt, and a GMV of [GMV]. It ranks within the top [sales volume ranking in leaf category] % in sales volume in the leaf category, with an average transaction value of [average transaction value]. Historical Time-based Average Performance: This product has undergone continuous advertisement exposure for number of advertising days days. The average hourly advertising spend distribution per day (from 0:$0$ to 24:$0$) is historical spend distribution sequence. The average GMV distribution across its category during this period is historical GMV distribution sequence. The average daily advertising budget is daily advertising budget.
D.1.2 Pair-wise Loss
Unlike human feedback scores used in LLM post-training (Christiano et al., 2017)—which can suffer from subjective biases in their absolute valuations—the trajectory quality $y(\tau)$ has real physical meaning and is comparable among different trajectories. Therefore, we can adopt a hybrid point-wise and pair-wise loss for the score to capture the absolute value of $y(\tau)$ and their relative preference, respectively. This approach has demonstrated superior performance in the recommendation domain (Cao et al., 2007; Lei et al., 2017; Wang et al., 2022a). Specifically, the pair-wise loss function can be expressed as:
$\mathbb{E}_{(\tau_{w},\tau_{l})\sim\mathcal{D}_{p}}\bigg[\log\sigma(\hat{y}_{\phi}(\tau_{w})-\hat{y}_{\phi}(\tau_{l}))\bigg],$ (50)
where the pair-wise loss is implemented by the typical Bradley-Terry (BT) function (Bradley & Terry, 1952). Here, $\mathcal{D}_{p}=\{(\tau_{w},\tau_{l},y(\tau_{w}),y(\tau_{l}))\}$ denotes a pair-wise dataset extracted from the offline dataset $\mathcal{D}$, where $\tau_{w}$ denotes the trajectory with higher trajectory quality, i.e., $y(\tau_{w})\geq y(\tau_{l})$.
D.1.3 Overall Evaluator Loss
Combining the previous techniques, the overall evaluator loss is:
$l_{e}(\phi)$ $=\underbrace{\mathbb{E}_{\tau\sim\mathcal{D}}\bigg[(\hat{y}_{\phi}(\tau)-y(\tau))^{2}\bigg]}_{\text{point-wise loss}}+\beta_{4}\underbrace{\mathbb{E}_{(\tau_{w},\tau_{l})\sim\mathcal{D}_{p}}\bigg[\log\sigma(\hat{y}_{\phi}(\tau_{w})-\hat{y}_{\phi}(\tau_{l}))\bigg]}_{\text{pair-wise loss}}$
$+\beta_{1}\underbrace{\mathbb{E}_{\tau_{1},\tau_{2}}\bigg[|\hat{y}_{\phi}(\tau_{1})-\hat{y}_{\phi}(\tau_{2})|-\sqrt{T}R_{m}\left\|\tau_{1}-\tau_{2}\right\|_{F}\bigg]_{+}}_{\text{Lipschitz penalty}},$ (51)
where $\beta_{4}>0$ is a hyper-parameter.
Figure 6: AIGB-Pearl Evaluator Learning. The evaluator loss is composed of three parts, including the point-wise loss, the pair-wise loss, and the Lipschitz penalty. The input representation of the evaluator is augmented with the LLM embedding to incorporate semantic information from textual advertiser features.
D.2 Algorithm Summary
Algorithm 1 summarizes the training process of AIGB-Pearl. Specifically, we calculate the Lipschitz constants of the $y(\tau)$ and $p_{D}(\tau|y)$ using the offline dataset $\mathcal{D}$, and accordingly, determine the Lipschitz constraint values for the evaluator and the planner, respectively. Then, we consecutively perform evaluator learning, planner pre-training, and KL-Lipschitz constrained score maximization for the planner. The development of AIGB-Pearl is supported by ROLL (Wang et al., 2025).
@@ D.2 Algorithm Summary
Input : Offline dataset $\mathcal{D}$, desired condition $y^{*}$, hyper-parameters $\beta_{1}$, $\beta_{2},\beta_{3},\beta_{4}$.
@@
// Determine the Lipschitz Constant
@@
Set the Lipschitz constraint value $\sqrt{T}R_{m}$ for the evaluator and $L_{p}$ for the planner, respectively.
// Learn the Trajectory Evaluator
@@
// Pre-Train the Generative Planner
while not converged do
  Update $\theta$ by minimizing Eq. 3;
end while
// KL-Lipschitz Constrained Score Maximization
@@
  Update $\theta$ by minimizing Eq. 3.2.2.
@@ E.3 Real-world Experiments on TargetROAS Bidding Problem
Table 7: The overall performance of AIGB-Pearl in TargetROAS problem in real-world A/B test, involving 300k advertisers over 22 days.
@@ E.4 Pathological Trajectory Behavior Explanation
E.5 Planner Training Stability
We present additional comparisons between the planner’s training curves of the offline RL with bootstrapping and those of AIGB-Pearl in Fig. 8, Fig. 9, Fig. 10, and Fig. 11 concerning:
@@ E.5 Planner Training Stability
We can observe that the offline RL method tends to exhibit significant instability throughout training, displaying high variance across different seeds. In contrast, AIGB-Pearl achieves much smoother and more consistent learning progress, demonstrating the improved training stability.
E.6 Evaluator Accuracy
Accuracy Metric. We evaluate the accuracy of the evaluator along two dimensions, including the absolute accuracy, reflecting how close the predicted scores are to ground truth qualities, and the order accuracy, reflecting the correctness of relative rankings between trajectory pairs. Specifically, we use the symmetric mean absolute percentage error, SMAPE, as the metric for the absolute accuracy and the AUC, defined as the ratio of correctly predicted ordinal pairs to the total number of pairs, as the metric for the order accuracy. The SMAPE ranges from $0\%$ to $200\%$ and AUC from $0\%$ to $100\%$. Lower SMAPE and larger AUC indicate better evaluator accuracy.
In the following, we investigate the effectiveness of employing LLM embedding and pair-wise loss in the evaluator learning.
LLM Embedding Effectiveness. We examine the accuracy of the trajectory evaluator without using LLM embeddings, and the results are reported in the lines of “w/o LLM” in Table 9 and Table 9. We can see that LLM embeddings can help improve both absolute and order accuracy. Fig. 7 compares the training progress with and without LLM embeddings in terms of the SMAPE metric. As illustrated, the evaluator incorporating LLM embeddings exhibits faster convergence and achieves lower SMAPE compared to the one without LLM embeddings. The performance gain stems from the ability of LLM embeddings to encode high-level semantic information, facilitating a more nuanced understanding of the dependencies between sequential bidding states.
Figure 7: The SMAPE training curves of the trajectory evaluator with and without LLM Embeddings. Incorporating LLM embeddings helps to achieve faster convergence and improved absolute accuracy.
Hybrid Point-wise and Pair-wise Losses Effectiveness. Table 9 and Table 9 present the SMAPE and AUC of the trajectory evaluator when trained with point-wise loss only, pair-wise loss only, and a combination of both. It can also be seen that using only pair-wise loss results in significantly worse SMAPE performance despite some improvement in AUC. This suggests that while pair-wise loss can enhance ranking consistency, it falls short in providing accurate absolute value predictions. When both point-wise and pair-wise losses are used together, the evaluator achieves lower SMAPE and higher AUC. This indicates that combining these two types of losses not only improves the absolute accuracy but also enhances the order accuracy in trajectory evaluation.
Table 8: SMAPE results from ablation experiments on the trajectory evaluator.
SMAPE $\downarrow$ Point-wise Pair-wise Both
w/o LLM 43.55% 196.38% 35.00%
with LLM 38.20% 196.00% 31.60%
Table 9: AUC results from ablation experiments on the trajectory evaluator.
AUC $\uparrow$ Point-wise Pair-wise Both
w/o LLM 61.57% 64.91% 70.91%
with LLM 65.00% 71.30% 75.10%